On the blueprint of a house, 27 millimeters represents 8 meters. The actual length of the living room is 10 meters. What is its length on the blueprint?

Answer by Theo(13342)

10/8 * 27 = 33.75 millimeters

10 is 1.25 times 8.

since 8 = 27 millimeters, and 10 = 1.25 times 8, then 1.25 times 27 millimeters = 33.75 millimeters.

another way of looking at it is:

27 millimeters = 8 meters.

27/8 = 3.375 millimeters on the blue print per 1 meter in real life.

10 meters in real life = 10 * 3.375 millimters per meter = 33.75 millimeters on the blueprint.